OpenStudy (anonymous):

Probability: I need help. So there are 6 colored balls and 3 are black. Ramona picks on and keeps it. Carlos picks one up too. What is the probability they with both pick a black ball. So its 3/6. Ramona has a 1/2 chance to pick a black ball. But then when she keeps it Carlos has a 2/5 chance to pick a black ball. I think im supposed to multiply but Im not too sure.

OpenStudy (kropot72):

The two events are independent, therefore the two probabilities should be multiplied.

jimthompson5910 (jim_thompson5910):

It's not independent since "Ramona picks one and keeps it. "

jimthompson5910 (jim_thompson5910):

there's no replacement

OpenStudy (anonymous):

So what do I do? multiply?

jimthompson5910 (jim_thompson5910):

you still multiply though P(both pick black) = P(Ramona picks black)*P(Carlos picks black based off Ramonas pick) P(both pick black) = (3/6)*(2/5) P(both pick black) = (3*2)/(6*5) P(both pick black) = 6/30 P(both pick black) = 1/5 This is assuming no replacement is made after Ramona picks a ball

OpenStudy (kropot72):

The events are independent, the reason being that the occurrence of one has absolutely no effect on the occurrence of the other. @jim_thompson5910 you are confusing occurrence with probability.

OpenStudy (anonymous):

thank you @jim_thompson5910 !! you're amazing :)

jimthompson5910 (jim_thompson5910):

No the events cannot be independent What Ramona picks changes what Carlos will pick

jimthompson5910 (jim_thompson5910):

they can only be independent if the ball Ramona picks is replaced or put back

OpenStudy (kropot72):

Sorry @jim_thompson5910 the events are independent and the correct result is obtained when the two probabilities found by @airrbae are multiplied.

jimthompson5910 (jim_thompson5910):

yes multiplication does occur, but they're still not independent if one event affects the probability of the other, then one event is dependent on the other (and they are not independent)

OpenStudy (kropot72):

When the probability is calculated using the hypergeometric distribution (sampling without replacement) the value obtained is consistent with the proof of independence which is \[P(A\cap B)=P(A)\times P(B)\]

jimthompson5910 (jim_thompson5910):

The formula used would be P(A and B) = P(A)*P(B | A) Now if A and B were independent, then P(B | A) = P(B) since event A wouldn't change P(B) at all which is why you can only say P(A and B) = P(A)*P(B) if A and B were independent events

jimthompson5910 (jim_thompson5910):

but they are not independent events Carlos has 1 less ball to choose from when Ramona picks out her ball, so the probability will be altered

OpenStudy (kropot72):

Whether or not Ramona has picked a ball has no effect on whether or not Carlos picks a ball. Independence relates to the occurrence of an event (picking a ball) and not to whether the probability of a certain outcome from an event (picking a ball) is affected by the outcome of a previous event. In any case the proof of independence that I gave should settle the matter.

jimthompson5910 (jim_thompson5910):

you have to define your events carefully I'm talking about the event of picking a black ball you're just talking about picking any ball in general

jimthompson5910 (jim_thompson5910):

we're focused on the probability of picking a black ball for both people so that's why A = event of Ramona picking a black ball B = event of Carlos picking a black ball

OpenStudy (kropot72):

@jim_thompson5910 Thank you very much. I accept your explanation that conditional probability is the key to understanding the correct theory. Thank you for your help :)

jimthompson5910 (jim_thompson5910):

you're welcome